I'm supposed to find out what x is in this problem:
2x+9y=-32
-5x-9y=26
I don't know how to get x by itself help!

Answer by checkley77(12569)

2x+9y=-32
-5x-9y=26 ADD.
--------------
-3X=-6
X=-6/-3
X=2 ANS.
2*2+9Y=-32
4+9Y=-32
9Y=-32-4
9Y=-36
Y=-36/9
Y=-4 ANS.
PROOF:
-5*2-9*-4=26
-10+36=26
26=26